A sample of hydrogen is collected at 34°C and heated to 68°C. At 68°C the volume is 26.0 L. What was the volume at 34°C?
Answer #1


using the charles law at constant pressure volume and temperature relation ship i.e volume is directly proportional to temperature is as follows

V1 / T1 = V2 / T2

here V1 = 26.0 L , T1 = 68 c = 341 K

         V2 = ? , T2 = 34 c = 307 K

substituting the values in the equation

V2 = V1 * T2 / T1

V2 = 26 * 307 / 341

V2 = 23.4076 L

answer => 23.4 L